Leon is four times as old as Matthew. Ten years from now, Leon will be twice as old as Matthew. Find the present ages of Leon and Matthew.
Answer by TimothyLamb(4379)

---
x = Leon
y = Mathew
---
x = 4y
x + 10 = 2(y + 10)
---
put the system of linear equations into standard form:
---
x - 4y = 0
x + 10 = 2y + 20
---
x - 4y = 0
x - 2y = 10
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = Leon = 20
y = Mathew = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php